40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1594     Entry time: Tue Oct 20 19:43:05 2015
Author: Antonio/Aidan/Eric 
Type: DailyProgress 
Category: optic 
Subject: Optics inventory 

The work that we need to do for the TCN experiment is at the state where we cannot easily go further without improving the lab organization.

I have collected the optics (1064nm) and made an inventory which is on dokuwiki: https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:experiments:psl:optics. At moment is little but I hope we can keep this updated and we can follow the policy that is written there.

The inventory is shown in the picture.

  • Yesterday and today some cleaning has been done by Eric and Aidan, removing big useless parts. The rack on the left is free now.
ELOG V3.1.3-